OpenStudy (anonymous):

Which of the following is an identity? A. sinx - cosx + 1 = tanx B. (1 - 2sin2x)csc2x = 4cos2x - 2 C. tan2x + cot2x = 1 D. sin2x cot2x + cos2x tan2x = 1

OpenStudy (jhannybean):

A looks like it has stuff added to it. \[\large \sin^2(x)+\cos^2(x) = 1\]

OpenStudy (anonymous):

what do you mean?

OpenStudy (jhannybean):

We know the identity s^2 +c^2 = 1. But whats tan(x) doing there?

OpenStudy (anonymous):

idk :c lol

OpenStudy (anonymous):

it shouldn't be there

OpenStudy (anonymous):

so the answer is A ?

OpenStudy (jhannybean):

Nope, i'm just going down the list one by one :D

OpenStudy (anonymous):

lol c: i think its B though like my gut is leaning towards it :D

OpenStudy (jhannybean):

Lets go with B now. \[\large (1-2\sin^2(x))(\csc^2(x))\]\[\large \csc^2(x) - 2\sin^2(x)\csc^2(x)\]hmmm...

OpenStudy (anonymous):

idk it could be C

jimthompson5910 (jim_thompson5910):

it's actually D sin^2x cot^2x + cos^2x tan^2x = 1 sin^2x (cos^2x/sin^2x) + cos^2x (sin^2x/cos^2x) = 1 cos^2x + sin^2x = 1 1 = 1

OpenStudy (anonymous):

@jim_thompson5910 thank you !!! :)

jimthompson5910 (jim_thompson5910):

np

OpenStudy (jhannybean):

I'm still relearning trig too. haha.